A random sample of Ersatz University students revealed that 10 females had a mean of $22.30 in their wallets with a standard deviation of $3.20, while 6 males had a mean of $17.30 with a standard deviation of $9.60. The critical value of F for an upper-tailed F-test for equal variances is:

a. 4.06

b. 0.287

c. 4.77

d. 3.48
Answer #1

since males have higher standard deviaiton, we wll treat that as first sample.

n1    = 6
n2    = 10
df of numerator v1 =n1-1=      5
df of denominator v2=n2-1 = 9
for 5 % level and given df critical values of F      =3.48

option D is correct